OpenStudy (anonymous):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is _____. the options -5 -1 1/2 1

OpenStudy (tester97):

I started solving this problem and realized i did it wrong. Maybe @dan815 Can help you?

OpenStudy (shinalcantara):

5x + y = 6 ---equation 1 5x + 3y = -4 ----equation 2 from equation 1 5x + y = 6 y = -5x + 6 ----equation 1' substitute equation 1' to equation 2 5x + 3y = -4 5x + 3(-5x+6) = -4 5x -15x + 18 = -4 5x - 15x = -4 - 18 -10x = -22 x = 22/10 x = 11/5 substitute x=11/5 to equation 1' y = -5x + 6 y = -5(11/5) + 6 y = -11 + 6 y = -5

OpenStudy (dan815):